# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ G.1 Motivation for tree-based wavelet filtering
To refine the hierarchical representations learned through our iterative learning scheme, we incorporate a filtering step $\Psi(\cdot)$ at each iteration. This filtering is based on Haar wavelets [Mallat, 1989, 1999, Haar, 1911, Daubechies, 1990] induced by the learned trees. Each sample ${\mathbf{X}}_{i,:}^{\top}\in\mathbb{R}^{m}$ can be viewed as a signal defined over the feature tree $\mathcal{T}(\mathbf{M}_{c})$, whose $m$ leaves correspond to features [Ortega et al., 2018]. Given a feature tree $\mathcal{T}(\mathbf{M}_{c})$, we construct an orthonormal Haar basis $\mathbf{B}_{c}\in\mathbb{R}^{m\times m}$, as described in Sec. 3 and App. A. Each sample is expanded in this basis, yielding coefficients $\bm{\alpha}_{i}=({\mathbf{X}}_{i,:}\mathbf{B}_{c})^{\top}\in\mathbb{R}^{m}$, where $\bm{\alpha}_{i}(j)=\langle{\mathbf{X}}_{i,:}^{\top},\bm{\beta}_{j}\rangle$ and $\bm{\beta}_{j}$ is the $j$-th wavelet in $\mathbf{B}_{c}$. To define a wavelet filter, we select a subset of the Haar basis vectors as follows. For each coefficient index $j$, we compute the aggregate $L_{1}$ norm $\sum_{i=1}^{n}|\bm{\alpha}_{i}(j)|$ across samples and sort the indices in descending order. We then select the subset $\Omega$ such that the cumulative contribution $\eta_{\Omega}=\sum_{q\in\Omega}\sum_{i=1}^{n}|\bm{\alpha}_{i}(q)|$ exceeds a threshold $\vartheta_{c}>0$. A similar filtering procedure can be applied to features using the Haar wavelet construct by the sample tree.
It is important to note that this $L_{1}$-based filtering retains the components that contribute most to the hierarchical decomposition, as represented by the tree [Coifman and Wickerhauser, 1992]. We postulate that the residual, i.e., the remaining coefficients outside the set $\Omega$, is the variation that is not aligned with the tree structure. These components are treated as noise or nuisance terms that are irrelevant for the hierarchical representation learning. By discarding this residual, the filter emphasizes signal content that is structurally consistent with the hierarchical structure, leading to representations that are more stable and meaningful across iterations [Grossmann and Morlet, 1984, Ngo et al., 2023]. Our filtering step adapts to the learned trees at each iteration. That is, the Haar wavelets are constructed from the trees $\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)})$ and $\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)})$ via TWD, making the filtering aligned with the evolving hierarchical representations. Empirically, we show that applying the wavelet filtering in the iterative learning scheme improves performance in various tasks on hierarchical data, including link prediction and node classification on hierarchical graph data, as well as sparse approximation and metric learning on scRNA-seq and document datasets.
@@ G.2 Advantages of Haar wavelet filtering over Laplacian-based filtering for hierarchical data
In Sec. 4.2, we introduced a filtering step based on the Haar wavelet, where the wavelet basis is induced by the tree structure [Gavish et al., 2010] learned during our iterative scheme. This filtering step is designed to reflect the intrinsic hierarchical structure of the data, enhancing meaningful hierarchical representations while suppressing noise and other nuisance components. An alternative filtering approach could involve using the Laplacian matrix derived from the tree graph. This would allow spectral filtering based on the eigendecomposition of the tree [Ortega et al., 2018, Dong et al., 2020], offering another way to apply filtering that incorporates the hierarchical information from the tree [Grone et al., 1990].
Consider a feature tree $\mathcal{T}(\mathbf{M}_{c})$ with the affinity matrix $\mathbf{A}_{c}\in\mathbb{R}^{m\times m}$. We can construct the Laplacian matrix for the feature tree by $\mathbf{L}_{c}=\mathbf{D}_{c}-\mathbf{A}_{c}$, where $\mathbf{D}_{c}$ is a diagonal degree matrix with $\mathbf{D}_{c}(j,j)=\sum_{i}\mathbf{A}(i,j)$. The eigenvectors of the Laplacian matrix are then treated as an orthonormal basis, often referred to as the graph Fourier basis [Chung, 1997, Stoica et al., 2005, Sandryhaila and Moura, 2013]. Subsequently, each sample can be expanded in this eigenbasis, and we denote $\widetilde{\bm{\alpha}}_{i}=({\mathbf{X}}_{i,:}\mathbf{U}_{c})^{\top}\in\mathbb{R}^{m}$ as the vector of the expansion coefficients of the $i$-th sample, where $\mathbf{U}_{c}$ is the matrix consisting of eigenvectors of $\mathbf{L}_{c}$. This expansion allows the interpretation of the Laplacian expansion coefficient $\widetilde{\bm{\alpha}}_{i}(j)$ as the contribution of the $j$-th frequency component to the signal. By sorting these coefficients according to the corresponding Laplacian eigenvalues, we can construct a low-pass filter (containing the first few expansions) that suppresses high-frequency components and noise while promoting signal smoothness [Hu et al., 2021]. A similar decomposition can also be applied to the features using the Laplacian matrix of the sample tree.
@@
We attribute this performance gap to the differences between the two filtering methods. Laplacian-based filtering is not designed to represent hierarchical structure explicitly [Ortega et al., 2018, Shuman et al., 2013]. Its eigenvectors are global and do not distinguish between different levels of abstraction in the data. In addition, from a computational perspective, computing the eigendecomposition of the graph Laplacian can be expensive: full decomposition scales cubically with the number of nodes, while partial decomposition (e.g., top $k$ eigenvectors) still incurs significant cost, often quadratic or worse in $n$ (depending on the method and graph sparsity). Furthermore, it has been shown that the expansion coefficients in the Haar wavelet basis decay at an exponential rate (Prop. A.1), in contrast to the typically polynomial rate of decay observed with Laplacian eigenfunction coefficients [Gavish et al., 2010]. This rapid decay reflects the ability of Haar wavelets to provide compact representations of piecewise-smooth signals, using the learned tree structure to concentrate energy in fewer coefficients. Finally, Haar wavelets on trees yield interpretable components [Mallat, 1989]: each wavelet coefficient corresponds to a specific scale and position within the tree, often representing differences between sibling nodes or between a node and its parent. This locality and hierarchical structure are absent in Laplacian eigenbases. Therefore, for data where a hierarchical relationship between samples or features is important, Haar wavelets provide a more structurally aligned approach compared to graph signal filtering based on the Laplacian.
We note that more advanced filtering strategies, such as those learned through neural networks like GCNs [Kipf and Welling, 2016], could also be explored. We leave this extension for future work.
@@ G.3 Comparison with Tree-WSV and the role of hierarchical representation learning
While preparing this manuscript, a new method called Tree-WSV [Düsterwald et al., 2025] was introduced. It uses Tree-Wasserstein Distance (TWD) to place samples and features as leaves on trees and was introduced to reduce the complexity that follows ideas from existing TWD research [Le et al., 2019, Yamada et al., 2022]. Specifically, this approach speeds up WSV [Huizing et al., 2022] for unsupervised ground metric learning.
The main idea of WSV is to simultaneously compute the Wasserstein distance among samples (resp. features) by using the Wasserstein distance among features (resp. samples) as the ground metric. This design conceptually aligns with co-manifold learning methods, which use relationships among samples to guide the relationships among features [Ankenman, 2014, Gavish and Coifman, 2012, Yair et al., 2017, Leeb and Coifman, 2016, Mishne et al., 2016, Shahid et al., 2016, Mishne et al., 2019, Chi et al., 2017]. Tree-WSV extends WSV by using TWD as a low-rank approximation of the Wasserstein distance. It does this by computing TWD among samples (resp. features) while using TWD among features (resp. samples) as the ground metric. Tree-WSV reduces the complexity of WSV from $O(n^{2}m^{2}(n\log n+m\log m))$ to $O(n^{3}+m^{3}+mn)$, making it more efficient in practice. In this context, the constructed trees primarily serve as computational tools for approximating Wasserstein distances.
In contrast, our work takes a geometric approach where the trees induced by TWD are used to represent hierarchical structures among samples and features. Based on this geometric perspective, we propose Alg. 1 and Alg. 2, which jointly learn hierarchical representations for both samples and features. Rather than using TWD merely as a computationally efficient approximation, we treat it as an input distance to construct diffusion operators that encode manifold information [Kileel et al., 2021]. These operators are then used to construct trees based on hyperbolic embeddings and diffusion densities [Lin et al., 2025], which recursively guide the computation of TWD for the other data mode.
This geometric view relates more closely to co-manifold learning in spirit, which use the geometry of one data mode to inform the geometry of the other, and was not explored in the WSV or Tree-WSV frameworks. Our primary goal is not unsupervised ground metric learning (though this arises naturally) but rather joint samples and features hierarchical representations. We also use multiple resolutions to better represent the hierarchical structure of the data, similar to ideas found in tree-based Earth mover’s distance [Ankenman, 2014, Mishne et al., 2019].

### v3

@@ G.1 Motivation for tree-based wavelet filtering
To refine the hierarchical representations learned through our iterative learning scheme, we incorporate a filtering step $\Psi(\cdot)$ at each iteration. This filtering is based on Haar wavelets [38, 39, 40, 41] induced by the learned trees. Each sample ${\mathbf{X}}_{i,:}^{\top}\in\mathbb{R}^{m}$ can be viewed as a signal defined over the feature tree $\mathcal{T}(\mathbf{M}_{c})$, whose $m$ leaves correspond to features [42]. Given a feature tree $\mathcal{T}(\mathbf{M}_{c})$, we construct an orthonormal Haar basis $\mathbf{B}_{c}\in\mathbb{R}^{m\times m}$, as described in Sec. 3 and App. A. Each sample is expanded in this basis, yielding coefficients $\bm{\alpha}_{i}=({\mathbf{X}}_{i,:}\mathbf{B}_{c})^{\top}\in\mathbb{R}^{m}$, where $\bm{\alpha}_{i}(j)=\langle{\mathbf{X}}_{i,:}^{\top},\bm{\beta}_{j}\rangle$ and $\bm{\beta}_{j}$ is the $j$-th wavelet in $\mathbf{B}_{c}$. To define a wavelet filter, we select a subset of the Haar basis vectors as follows. For each coefficient index $j$, we compute the aggregate $L_{1}$ norm $\sum_{i=1}^{n}|\bm{\alpha}_{i}(j)|$ across samples and sort the indices in descending order. We then select the subset $\Omega$ such that the cumulative contribution $\eta_{\Omega}=\sum_{q\in\Omega}\sum_{i=1}^{n}|\bm{\alpha}_{i}(q)|$ exceeds a threshold $\vartheta_{c}>0$. A similar filtering procedure can be applied to features using the Haar wavelet construct by the sample tree.
It is important to note that this $L_{1}$-based filtering retains the components that contribute most to the hierarchical decomposition, as represented by the tree [69]. We postulate that the residual, i.e., the remaining coefficients outside the set $\Omega$, is the variation that is not aligned with the tree structure. These components are treated as noise or nuisance terms that are irrelevant for the hierarchical representation learning. By discarding this residual, the filter emphasizes signal content that is structurally consistent with the hierarchical structure, leading to representations that are more stable and meaningful across iterations [166, 167]. Our filtering step adapts to the learned trees at each iteration. That is, the Haar wavelets are constructed from the trees $\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)})$ and $\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)})$ via TWD, making the filtering aligned with the evolving hierarchical representations. Empirically, we show that applying the wavelet filtering in the iterative learning scheme improves performance in various tasks on hierarchical data, including link prediction and node classification on hierarchical graph data, as well as sparse approximation and metric learning on scRNA-seq and document datasets.
@@ G.2 Advantages of Haar wavelet filtering over Laplacian-based filtering for hierarchical data
In Sec. 4.2, we introduced a filtering step based on the Haar wavelet, where the wavelet basis is induced by the tree structure [43] learned during our iterative scheme. This filtering step is designed to reflect the intrinsic hierarchical structure of the data, enhancing meaningful hierarchical representations while suppressing noise and other nuisance components. An alternative filtering approach could involve using the Laplacian matrix derived from the tree graph. This would allow spectral filtering based on the eigendecomposition of the tree [42, 168], offering another way to apply filtering that incorporates the hierarchical information from the tree [169].
Consider a feature tree $\mathcal{T}(\mathbf{M}_{c})$ with the affinity matrix $\mathbf{A}_{c}\in\mathbb{R}^{m\times m}$. We can construct the Laplacian matrix for the feature tree by $\mathbf{L}_{c}=\mathbf{D}_{c}-\mathbf{A}_{c}$, where $\mathbf{D}_{c}$ is a diagonal degree matrix with $\mathbf{D}_{c}(j,j)=\sum_{i}\mathbf{A}(i,j)$. The eigenvectors of the Laplacian matrix are then treated as an orthonormal basis, often referred to as the graph Fourier basis [170, 171, 172]. Subsequently, each sample can be expanded in this eigenbasis, and we denote $\widetilde{\bm{\alpha}}_{i}=({\mathbf{X}}_{i,:}\mathbf{U}_{c})^{\top}\in\mathbb{R}^{m}$ as the vector of the expansion coefficients of the $i$-th sample, where $\mathbf{U}_{c}$ is the matrix consisting of eigenvectors of $\mathbf{L}_{c}$. This expansion allows the interpretation of the Laplacian expansion coefficient $\widetilde{\bm{\alpha}}_{i}(j)$ as the contribution of the $j$-th frequency component to the signal. By sorting these coefficients according to the corresponding Laplacian eigenvalues, we can construct a low-pass filter (containing the first few expansions) that suppresses high-frequency components and noise while promoting signal smoothness [173]. A similar decomposition can also be applied to the features using the Laplacian matrix of the sample tree.
@@
We attribute this performance gap to the differences between the two filtering methods. Laplacian-based filtering is not designed to represent hierarchical structure explicitly [42, 72]. Its eigenvectors are global and do not distinguish between different levels of abstraction in the data. In addition, from a computational perspective, computing the eigendecomposition of the graph Laplacian can be expensive: full decomposition scales cubically with the number of nodes, while partial decomposition (e.g., top $k$ eigenvectors) still incurs significant cost, often quadratic or worse in $n$ (depending on the method and graph sparsity). Furthermore, it has been shown that the expansion coefficients in the Haar wavelet basis decay at an exponential rate (Prop. A.1), in contrast to the typically polynomial rate of decay observed with Laplacian eigenfunction coefficients [43]. This rapid decay reflects the ability of Haar wavelets to provide compact representations of piecewise-smooth signals, using the learned tree structure to concentrate energy in fewer coefficients. Finally, Haar wavelets on trees yield interpretable components [38]: each wavelet coefficient corresponds to a specific scale and position within the tree, often representing differences between sibling nodes or between a node and its parent. This locality and hierarchical structure are absent in Laplacian eigenbases. Therefore, for data where a hierarchical relationship between samples or features is important, Haar wavelets provide a more structurally aligned approach compared to graph signal filtering based on the Laplacian.
We note that more advanced filtering strategies, such as those learned through neural networks [96], could also be explored. We leave this extension for future work.
@@ G.3 Comparison with Tree-WSV and the role of hierarchical representation learning
While preparing this manuscript, a new method called Tree-WSV [60] was introduced. It uses Tree-Wasserstein Distance (TWD) to place samples and features as leaves on trees and was introduced to reduce the complexity that follows ideas from existing TWD research [35, 34]. Specifically, this approach speeds up WSV [59] for unsupervised ground metric learning.
The main idea of WSV is to simultaneously compute the Wasserstein distance among samples (resp. features) by using the Wasserstein distance among features (resp. samples) as the ground metric. This design conceptually aligns with co-manifold learning methods, which use relationships among samples to guide the relationships among features [52, 53, 54, 55, 7, 56, 57, 63]. Tree-WSV extends WSV by using TWD as a low-rank approximation of the Wasserstein distance. It does this by computing TWD among samples (resp. features) while using TWD among features (resp. samples) as the ground metric. Tree-WSV reduces the complexity of WSV from $O(n^{2}m^{2}(n\log n+m\log m))$ to $O(n^{3}+m^{3}+mn)$, making it more efficient in practice. In this context, the constructed trees primarily serve as computational tools for approximating Wasserstein distances.
In contrast, our work takes a geometric approach where the trees induced by TWD are used to represent hierarchical structures among samples and features. Based on this geometric perspective, we propose Alg. 1 and Alg. 2, which jointly learn hierarchical representations for both samples and features. Rather than using TWD merely as a computationally efficient approximation, we treat it as an input distance to construct diffusion operators that encode manifold information [68]. These operators are then used to construct trees based on hyperbolic embeddings and diffusion densities [37], which recursively guide the computation of TWD for the other data mode.
This geometric view relates more closely to co-manifold learning in spirit, which use the geometry of one data mode to inform the geometry of the other, and was not explored in the WSV or Tree-WSV frameworks. Our primary goal is not unsupervised ground metric learning (though this arises naturally) but rather joint samples and features hierarchical representations. We also use multiple resolutions to better represent the hierarchical structure of the data, similar to ideas found in tree-based Earth mover’s distance [52, 57].
@@
G.4 On the interpretability of TWDs
Note that one strength of TWD lies in its transparent and decomposable structure. Given two distributions (e.g., feature or sample vectors), the TWD quantifies the minimal “mass transport” needed to align them over the tree, with transport cost reflecting branch length. This formulation enables users to understand where and at what scale differences occur. For instance, if most of the transport cost is concentrated at high-level branches, the two distributions differ in broad terms; if the cost is concentrated near the leaves, they are largely similar but differ in fine details. As a result, TWD not only quantifies similarity but also supports qualitative interpretation: users can trace which parts of the tree contribute most to the distance between any two data points. In classification tasks, this interpretability reveals which regions of the hierarchy drive class separation and can help identify key transitions in the data.